Find the slope and y-intercept of the line represented by the following equation.
2x - 3y = 6

Answer by 303795(602)

Rearrange the equation to get y on its own (and positive).
2x - 3y = 6
Add 3y to each side of the equation to get a positive y.
2x - 3y +3y = 6 + 3y
2x = 6 + 3y
take 6 from each side of the equation.
2x - 6 = 6 + 3y - 6
2x - 6 = 3y
Divide each side of the equation by 3.
(2x - 6)/3 = 3y/3
2/3x - 2 = y
y = 2/3x - 2
So the slope is 2/3 and the y intercept is -2